A truck’s speed increases uniformly from 15 km/h to 60 km/h in 20 s. Determine (a) the average speed, (b) the acceleration, and (c) the distance travelled, all in units of meters and seconds.


1
Expert's answer
2021-06-21T11:39:11-0400

(a) The average speed:


"v=\frac{v_1+v_2}{2}=10.4\text{ m/s}."

(b) The acceleration:


"a=\frac{v_2-v_1}{t}=0.625\text{ m/s}^2."

(c) The distance is


"d=\frac{v_2^2-v_1^2}{2a}=208\text{ m}."